# Negate or Embrace: On How Misalignment Shapes Multimodal Representation Learning

### v3

@@ Abstract
Multimodal representation learning, exemplified by multimodal contrastive learning (MMCL) using image-text pairs, aims to learn powerful representations by aligning cues across modalities. This approach relies on the core assumption that the exemplar image-text pairs constitute two representations of an identical concept. However, recent research has revealed that real-world datasets often exhibit misalignment. There are two distinct viewpoints on how to address this issue: one suggests mitigating the misalignment, and the other leveraging it. We seek here to reconcile these seemingly opposing perspectives, and to provide a practical guide for practitioners. Using latent variable models we thus formalize misalignment by introducing two specific mechanisms: selection bias, where some semantic variables are missing, and perturbation bias, where semantic variables are distorted—both affecting latent variables shared across modalities. Our theoretical analysis demonstrates that, under mild assumptions, the representations learned by MMCL capture exactly the information related to the subset of the semantic variables invariant to selection and perturbation biases. This provides a unified perspective for understanding misalignment. Based on this, we further offer actionable insights into how misalignment should inform the design of real-world ML systems. We validate our theoretical findings through extensive empirical studies on both synthetic data and real image-text datasets, shedding light on the nuanced impact of misalignment on multimodal representation learning.
footnotetext: These authors contributed equally. Correspondence to yichao.cai@adelaide.edu.au.
@@ 1 Introduction
Modern multimodal learning has achieved remarkable success by jointly modeling information from heterogeneous sources such as vision, language, and audio. In particular, multimodal contrastive learning (MMCL) on paired data has emerged as a dominant strategy for aligning modalities (Radford et al., 2021; Jia et al., 2021; Wu et al., 2022)—notably exemplified by vision-language models like CLIP, which attempt to learns a common representation space by maximizing the similarity of real image-text pairs while minimizing that of incorrect pairs (Radford et al., 2021). However, one fundamental assumption in MMCL is that the training pairs are completely aligned across modalities, meaning that they convey exactly the same semantic information (Xia et al., 2023; Liu et al., 2024d). This assumption, though convenient, is often violated in real-world scenarios, where multimodal data is inherently noisy or imprecisely paired (Miech et al., 2019; Nakada et al., 2023), e.g., when the text only partially describes the paired image content, or even contains descriptive elements that are irrelevant or misleading, a situation generally referred to as misalignment. For example, in a large-scale video-text dataset, over 50% of the purportedly aligned clip-caption pairs were found to be misaligned (Miech et al., 2019). It is worth noting that the impact of misalignment is primarily concentrated in the text modality, as text captions—often generated based on images—are more susceptible to semantic incompleteness and interpretive variability compared to the typically fixed visual semantic aspects.
The misalignment discussed above has led to two seemingly opposing viewpoints. On one hand, misalignment is viewed as a form of disruption that should be mitigated (Liu et al., 2024a; Sun et al., 2024; Wu et al., 2022; Chen et al., 2024; Tao et al., 2024; Kim et al., 2024; Xuan et al., 2022; Chartsias et al., 2020; Lv et al., 2021). For example, misalignment between modalities can result in "hallucination" in multimodal models (Sun et al., 2024; Leng et al., 2024; Xie et al., 2024). It also provides weak, noisy, and even misleading supervision for multimodal pre-training (Wu et al., 2022; Zhang et al., 2024b). On the other hand, an opposing viewpoint suggests that multimodal representations may actually benefit from misalignment (Wu et al., 2023; Nakada et al., 2023; Cai et al., 2025; Kim et al., 2023). For instance, fine-tuning the representations learned by CLIP through random text augmentation—which deliberately introduces misalignment in style-related information—can lead to more robust representations for zero-shot learning, few-shot learning, and even adversarial attacks (Cai et al., 2025). This contrast raises a crucial question:
@@
In light of this, we offer a theoretical perspective that not only facilitates the understanding of misalignment but also provides insights into real-world applications. Specifically, we formulate the problem via a latent variable model (LVM), which depicts the underlying generative process of image-text data with misalignment, as shown in Figure 1. In it, the latent space consists of shared semantic variables representing factors common to both modalities (e.g., object shapes and colors), along with modality-specific subspaces that capture unique variations in images and text. To model misalignment, we introduce two mechanisms: selection bias and perturbation bias. Both act on the shared semantic information but differ in their effects. Selection bias determines which semantic information is preserved in the text. For example, when describing an object, the text might preserve information about its color ("black") but omit details about its texture or shape. On the other hand, perturbation bias introduces errors, such as changing "black" (correct color) to "red" (incorrect color). To make the proposed LVM adaptable to a wide range of real-world scenarios, we allow for an arbitrary causal structure among the latent semantic variables, providing flexibility in multimodal contexts. Finally, given their differences, we model the modalities with separate generative processes.
@@
Building on the proposed LVM (generative model), we present a theoretical identifiability analysis within MMCL framework (inference model). We show, under mild assumptions, that the subset of semantic variables unaffected by selection and perturbation biases remain block-identifiable (Defn. 4.1)—that is, only the unaffected subset of semantic variables admits a nonlinear and invertible mapping to the representations learned by MMCL in the proposed LVM. In contrast, the remaining semantic variables that are affected by misalignment are inherently excluded from the learned representations, regardless of the latent causal structure among all semantic variables. This result provides a unified perspective on the seemingly opposing views discussed above. While misalignment can be problematic in tasks that rely on fully preserving semantic information to maximize downstream utility, it may paradoxically become beneficial in scenarios where robustness to distribution shifts is desired. In such cases, misalignment acts as a natural regularizer, implicitly guiding models to focus on stable, invariant semantic factors shared across modalities.
@@
We summarize our key contributions as follows, and a detailed discussion of related work is provided in § 6. (i) We propose a latent variable model for multimodal data generation that explicitly captures misalignment through two mechanisms: selection bias and perturbation bias (§ 3). (ii) We establish a general identifiability result, showing that MMCL recovers the subset of semantic variables unaffected by these biases, independent of the underlying latent causal structure (§ 4.1). (iii) We extend this result to two practical scenarios—tasks requiring common representations and those targeting invariant representations—offering actionable insights into how misalignment should inform real-world applications (§ 4.2). (iv) We empirically validate our theoretical findings through extensive experiments on both real-world and synthetic image-text datasets, across diverse misalignment settings, including those with structured latent dependencies (§ 5).
@@ 2 Preliminaries: Multimodal Contrastive Learning
Multimodal contrastive learning (MMCL) (Radford et al., 2021; Jia et al., 2021; Zhang et al., 2022) aims to learn joint representations by aligning paired samples from different modalities, i.e., $\mathbf{t}\in\mathcal{T}$ for text and $\mathbf{x}\in\mathcal{X}$ for images, while pushing apart unpaired (negative) samples. In practice, MMCL typically employs two modality-specific encoders, i.e., $\mathbf{f}_{t}(\mathbf{t})$ for text and $\mathbf{f}_{x}(\mathbf{x})$ for images which project observed paired data into a shared representation space. The learning objective is generally formulated as the following contrastive loss:
$\mathcal{L}_{\text{MMCL}}(\mathbf{f}_{x},\mathbf{f}_{t})=-\frac{1}{2K}\left[\sum\limits_{i=1}^{K}\log\frac{e^{\langle\mathbf{f}_{x}(\mathbf{x}_{i}),\mathbf{f}_{t}(\mathbf{t}_{i})\rangle/\tau}}{\sum\nolimits_{j=1}^{K}e^{\langle\mathbf{f}_{x}(\mathbf{x}_{i}),\mathbf{f}_{t}(\mathbf{t}_{j})\rangle/\tau}}+\sum\limits_{i=1}^{K}\log\frac{e^{\langle\mathbf{f}_{x}(\mathbf{x}_{i}),\mathbf{f}_{t}(\mathbf{t}_{i})\rangle/\tau}}{\sum\nolimits_{j=1}^{K}e^{\langle\mathbf{f}_{x}(\mathbf{x}_{j}),\mathbf{f}_{t}(\mathbf{t}_{i})\rangle/\tau}}\right],$ (1)
@@
$\mathcal{L}_{\text{SymAlignMaxEnt}}(\mathbf{f}_{x},\mathbf{f}_{t})=\mathop{\mathbb{E}}_{(\mathbf{x},\mathbf{t})\sim p_{\mathbf{x},\mathbf{t}}}\left[\|\mathbf{f}_{x}(\mathbf{x})-\mathbf{f}_{t}(\mathbf{t})\|_{2}\right]-\frac{1}{2}\Bigl{(}H\bigl{(}\mathbf{f}_{x}(\mathbf{x})\bigr{)}+H\bigl{(}\mathbf{f}_{t}(\mathbf{t})\bigr{)}\Bigr{)},$ (2)
where $H(\cdot)$ denotes differential entropy (Wang and Isola, 2020; Von Kügelgen et al., 2021). One of the main advantages of the asymptotic objective in Eq. 2 is its suitability for theoretical analysis (Lyu et al., 2022; Daunhawer et al., 2022; Yao et al., 2023). At a high level, Eq. 2 naturally decomposes into two intuitive terms (Wang and Isola, 2020): the first term encourages minimizing the distance between paired samples, while the second term promotes maximizing the entropy of learned representations. Following prior works (Lyu et al., 2022; Daunhawer et al., 2022; Yao et al., 2023), we also adopt this objective in our theoretical analysis.
However, a key distinction of our work is its focus on the impact of misalignment, in contrast to prior studies that assume perfectly aligned paired data. To model this, we introduce a novel latent variable formulation (§ 3) that defines a fundamentally different problem setting. Consequently, our theoretical analysis (§ 4) yields insights that diverge significantly from existing results, highlighting how misalignment influences the learned representations in multimodal contrastive learning.
@@ 3 Problem Formulation via a Generative Perspective
In this section, we introduce a novel latent variable model (LVM) to formally characterize misalignment (§ 3.1). The model incorporates two key mechanisms, selection bias and perturbation bias, to explicitly capture distinct sources of misalignment. Building on this model, we present assumptions underlying image-text pairs under misalignment (§ 3.2). See App. A for a full notation table.
3.1 A Latent Variable Model Characterizing Misalignment
@@
  • Semantic variables $\mathbf{s}\in\mathcal{S}\subseteq\mathbb{R}^{n_{s}}$: Latent variables capturing the semantic content of the data, i.e., information that is interpretable or describable through human knowledge (e.g., object shape, color). We denote the index set of semantic variables as $\mathbb{I}_{\mathbf{s}}:=\{1,\dots,n_{s}\}$ for future reference.
  • Image-specific variables $\mathbf{m}_{x}\in\mathcal{M}_{x}$  [^1] For simplicity of notation, we omit the dimensions of certain variables throughout this work.: Latent variables capturing image-specific, non-semantic factors (e.g., camera noise or background artifacts) that are independent of semantic variables $\mathbf{s}$.
  • Text-specific variables $\mathbf{m}_{t}\in\mathcal{M}_{t}$: Latent variables capturing text-specific, non-semantic factors (e.g., grammar) that are independent of both semantic variables $\mathbf{s}$ and image-specific variables $\mathbf{m}_{x}$.
A key advantage of the proposed latent space structure is its flexibility and applicability to real-world scenarios. To this end, we depart from prior works that impose somewhat restrictive assumptions on the latent structure. Specifically, unlike approaches that enforce certain fixed graphical structures among semantic variables (e.g., assuming content causally determines style) (Daunhawer et al., 2022), or methods based on nonlinear ICA that assume complete (or conditional on an auxiliary variable) independence among latent variables (Khemakhem et al., 2020; Sorrenson et al., 2020), we allow for arbitrary dependency structures among semantic variables $\mathbf{s}$, since the true latent graph structure among semantic variables is often unknown in practice.
@@
Image $\mathbf{x}\in\mathcal{X}$ is generated from latent variables $\mathbf{z}_{x}=(\mathbf{s},\mathbf{m}_{x})$ via a diffeomorphism (i.e., a bijection with smooth inverse function) $\mathbf{g}_{x}:\mathcal{S}\times\mathcal{M}_{x}\to\mathcal{X}$:
@@
Unlike other multimodal data (e.g., camera-LiDAR (Liang et al., 2022)) acquired via sensors, given an image, its corresponding text inherently exhibits flexibility in semantic richness and may also include distortions introduced by humans or captioning models (Li et al., 2022; Xu et al., 2024), leading to misalignment. Here, we formalize such misalignment by introducing two types of biases: selection bias and perturbation bias. Accordingly, before formulating the text generation process, we first provide definitions of these two types of biases.
@@
Let $\mathcal{P}_{+}(\mathbb{I}_{\mathbf{s}})$ denote the set of all non-empty subsets [^2] Without loss of generality, we fix a graded lexicographic order over $\mathcal{P}_{+}(\mathbb{I}_{\mathbf{s}})$. See Clar. B.2 for a clarification. of the index set $\mathbb{I}_{\mathbf{s}}$, defined as $\mathcal{P}_{+}(\mathbb{I}_{\mathbf{s}}):=\mathcal{P}(\mathbb{I}_{\mathbf{s}})\setminus\{\emptyset\}$, where $\mathcal{P}(\cdot)$ denotes the power set. The selection bias $\theta$ is defined as an integer index in the range $\{1,\dots,2^{n_{s}}-1\}$, corresponding to a specific non-empty semantic subset $\mathbb{I}_{\theta}\in\mathcal{P}_{+}(\mathbb{I}_{\mathbf{s}})$. The complement $\mathbb{I}_{\theta}^{c}=\mathbb{I}_{\mathbf{s}}\setminus\mathbb{I}_{\theta}$ denotes the omitted semantic subset.
@@
Let $\mathcal{P}_{\text{proper}}(\mathbb{I}_{\theta}):=\mathcal{P}(\mathbb{I}_{\theta})\setminus\{\mathbb{I}_{\theta}\}$ denote the set of all proper subsets [^3] Again, we fix a graded lexicographic order over the proper subsets in $\mathcal{P}_{\text{proper}}(\mathbb{I}_{\theta})$ throughout the paper. of the selected index subset $\mathbb{I}_{\theta}$. The perturbation bias $\rho$ is defined as an integer index in the range $\rho\in\{1,\dots,2^{|\mathbb{I}_{\theta}|}-1\},$ corresponding to a unique subset $\mathbb{I}_{\rho}\in\mathcal{P}_{\text{proper}}(\mathbb{I}_{\theta})$ subject to perturbation. The complement $\mathbb{I}_{\rho}^{c}:=\mathbb{I}_{\theta}\setminus\mathbb{I}_{\rho}$ denotes the semantic components that remain fixed.
@@
where $p_{\mathbf{m}_{t}}$ denotes the prior distribution over the latent subspace $\mathcal{M}_{t}$, and $\mathbf{g}_{t^{\scriptscriptstyle(\theta)}}$ is the diffeomorphic mapping specified by the selection bias $\theta$. This formulation explicitly captures how misalignment in text generation arises through perturbations within selected semantic dimensions.

### v4

# On the Value of Cross-Modal Misalignment in Multimodal Representation Learning

@@ Abstract
Multimodal representation learning, exemplified by multimodal contrastive learning (MMCL) using image-text pairs, aims to learn powerful representations by aligning cues across modalities. This approach relies on the core assumption that the exemplar image-text pairs constitute two representations of an identical concept. However, recent research has revealed that real-world datasets often exhibit cross-modal misalignment. There are two distinct viewpoints on how to address this issue: one suggests mitigating the misalignment, and the other leveraging it. We seek here to reconcile these seemingly opposing perspectives, and to provide a practical guide for practitioners. Using latent variable models we thus formalize cross-modal misalignment by introducing two specific mechanisms: Selection bias, where some semantic variables are absent in the text, and perturbation bias, where semantic variables are altered—both leading to misalignment in data pairs. Our theoretical analysis demonstrates that, under mild assumptions, the representations learned by MMCL capture exactly the information related to the subset of the semantic variables invariant to selection and perturbation biases. This provides a unified perspective for understanding misalignment. Based on this, we further offer actionable insights into how misalignment should inform the design of real-world ML systems. We validate our theoretical findings via extensive empirical studies on both synthetic data and real image-text datasets, shedding light on the nuanced impact of cross-modal misalignment on multimodal representation learning.  [^1] The code is available at https://github.com/YichaoCai1/crossmodal_mislaignment
@@ 1 Introduction
Modern multimodal learning has achieved remarkable success in jointly modeling information from heterogeneous sources such as vision, language, and audio. Multimodal contrastive learning (MMCL) on paired data has emerged as a dominant strategy for aligning modalities [53, 25, 70]. This approach is exemplified by vision-and-language models like CLIP, which attempt to identify a common representation space that maximizes the similarity of real image-text pairs while minimizing that of incorrect pairs [53]. One of the assumptions underpinning this approach is that the training pairs are aligned across modalities, meaning that they convey exactly the same semantic information [71, 43]. This assumption, though convenient, is often violated in real-world scenarios, where multimodal data is inherently noisy or imprecisely paired [47, 48]. More critically, text taken from image captions typically only partially describes the paired image content, and often contains descriptive elements that are irrelevant or misleading, a situation referred to as cross-modal misalignment. For example, in a large-scale video-text dataset, over 50% of the purportedly aligned clip-caption pairs were found to be misaligned [47]. It is worth noting that the impact of cross-modal misalignment is mainly concentrated in the text modality, as text captions—often derived from images—are more susceptible to semantic incompleteness and interpretive variability than the typically consistent visual semantics.
The misalignment [^2] For the sake of brevity, we use“cross-modal misalignment” and “misalignment” interchangeably throughout the paper, as the context clearly indicates that we are referring to misalignment between different modalities. discussed above has led to two seemingly opposing viewpoints. On one hand, misalignment is viewed as a form of disruption that should be mitigated [38, 62, 70, 7, 63, 28, 74, 6, 45]. For example, cross-modal misalignment can result in “hallucination” in multimodal models [62, 33, 72]. It also provides weak, noisy, and even misleading supervision for multimodal pre-training [70, 79]. On the other hand, an alternative viewpoint suggests that multimodal representations may actually benefit from cross-modal misalignment [69, 48, 5, 27]. For instance, fine-tuning the representations learned by CLIP through random text augmentation—which deliberately introduces misalignment in style-related information—can lead to more robust representations for zero-shot learning, few-shot learning, and even adversarial attacks [5]. This contrast raises a crucial question:
@@
In light of this, we offer a theoretical perspective that not only facilitates the understanding of misalignment but also provides guidance for real-world applications. Specifically, we formulate the problem using a latent variable model (LVM), which depicts the underlying generative process of image-text data with misalignment, as shown in Figure 1. In it, the latent space consists of latent semantic variables representing factors common to both modalities (e.g., object shapes and colors), along with modality-specific subspaces that capture unique variations in images and text. To make the LVM more adaptable to diverse real-world scenarios, we allow for an arbitrary causal structure among the latent semantic variables, providing flexibility in multimodal contexts. To model misalignment, we introduce two mechanisms: selection bias and perturbation bias. Both act on latent semantic information but differ in their effects. Selection bias determines which semantic information is preserved in the text. For example, when describing an object, the text might preserve information about its color (“black”) but omit its texture details or shape. On the other hand, perturbation bias introduces errors, such as misannotating “black” (correct color) to “red” (incorrect color). Moreover, given their heterogeneity, we model the two modalities with separate generative processes.
@@
Building on the proposed LVM (i.e., the generative model), we present a theoretical identifiability analysis within the MMCL framework (i.e., the inference model). We show, under mild assumptions, that the subset of semantic variables unaffected by selection and perturbation biases remain block-identifiable (Defn. 4.1)—that is, only the unaffected subset of semantic variables in the proposed LVM admits an invertible mapping to the representations learned by MMCL. In contrast, the remaining misaligned semantic variables affected by either bias are inherently excluded from the learned representations, irrespective of the latent causal structure among all semantic variables. This result provides a unified perspective on the seemingly opposing views discussed above. While misalignment can be problematic in tasks that rely on fully preserving semantic information to maximize downstream utility, it may paradoxically become beneficial in scenarios where robustness to distribution shifts is desired. In such cases, cross-modal misalignment naturally acts as a regularizer, encouraging models to focus on stable, invariant semantic factors shared across modalities.
@@
Contributions. Our main contributions are: (i) We propose an LVM for multimodal data generation that explicitly characterizes cross-modal misalignment through two mechanisms: selection bias and perturbation bias (§ 3). (ii) We establish a general identifiability result, showing that MMCL recovers the subset of semantic variables unaffected by these biases, independent of the underlying latent causal structure (§ 4.1). (iii) We extend this result to two practical scenarios—tasks requiring common representations and those targeting invariant representations—offering actionable insights into how misalignment should inform real-world applications (§ 4.2). (iv) We empirically validate our theoretical findings through extensive experiments on both real-world and synthetic image-text datasets, across diverse cross-modal misalignment settings, including those with structured latent dependencies (§ 5). A comprehensive discussion of related work can be found in § 6.
@@ 2 Preliminaries: Multimodal Contrastive Learning
Multimodal contrastive learning (MMCL) [53, 25, 78] aims to learn joint representations by aligning paired samples from different modalities, i.e., $\mathbf{t}\in\mathcal{T}$ for text and $\mathbf{x}\in\mathcal{X}$ for images, while pushing apart unpaired (negative) samples. In practice, MMCL typically employs two modality-specific encoders, i.e., $\mathbf{f}_{t}(\mathbf{t})$ for text and $\mathbf{f}_{x}(\mathbf{x})$ for images which project observed paired data into a shared representation space. The learning objective is generally formulated as the following contrastive loss:
$\mathcal{L}_{\text{\tiny MMCL}}(\mathbf{f}_{x},\mathbf{f}_{t})=-\frac{1}{2K}\left[\sum\limits_{i=1}^{K}\log\frac{e^{\langle\mathbf{f}_{x}(\mathbf{x}_{i}),\mathbf{f}_{t}(\mathbf{t}_{i})\rangle/\tau}}{\sum\nolimits_{j=1}^{K}e^{\langle\mathbf{f}_{x}(\mathbf{x}_{i}),\mathbf{f}_{t}(\mathbf{t}_{j})\rangle/\tau}}+\sum\limits_{i=1}^{K}\log\frac{e^{\langle\mathbf{f}_{x}(\mathbf{x}_{i}),\mathbf{f}_{t}(\mathbf{t}_{i})\rangle/\tau}}{\sum\nolimits_{j=1}^{K}e^{\langle\mathbf{f}_{x}(\mathbf{x}_{j}),\mathbf{f}_{t}(\mathbf{t}_{i})\rangle/\tau}}\right],$ (1)
@@
$\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}(\mathbf{f}_{x},\mathbf{f}_{t})=\mathop{\mathbb{E}}\nolimits_{(\mathbf{x},\mathbf{t})\sim p_{\mathbf{x},\mathbf{t}}}\left[\|\mathbf{f}_{x}(\mathbf{x})-\mathbf{f}_{t}(\mathbf{t})\|_{2}\right]-\frac{1}{2}\Bigl{(}H\bigl{(}\mathbf{f}_{x}(\mathbf{x})\bigr{)}+H\bigl{(}\mathbf{f}_{t}(\mathbf{t})\bigr{)}\Bigr{)},$ (2)
where $H(\cdot)$ denotes differential entropy [67, 66]. One of the main advantages of the asymptotic objective in Eq. 2 is its suitability for theoretical analysis [46, 12, 76]. At a high level, Eq. 2 naturally decomposes into two intuitive terms [67]: the first term encourages minimizing the distance between paired samples, while the second term promotes maximizing the entropy of learned representations. Following prior works [46, 12, 76], we also adopt this objective in our theoretical analysis.
A key feature of our work is its focus on the impact of cross-modal misalignment, in contrast to prior studies that assume perfectly aligned paired data. To model this, we introduce a novel latent variable model (§ 3) that defines a fundamentally different problem setting. Consequently, our theoretical analysis (§ 4) yields insights that diverge significantly from existing results, highlighting how misalignment influences the learned representations in multimodal contrastive learning.
@@ 3 Problem Formulation via a Generative Perspective
In this section, we introduce a novel latent variable model (LVM) to formalize the underlying generative processes of image-text data, explicitly characterizing cross-modal misalignment (§ 3.1). Building on this LVM, we introduce the technical assumptions underlying image-text pairs under misalignment for subsequent analysis (§ 3.2). See App. A for a complete notation table.
3.1 A Latent Variable Model Characterizing Cross-Modal Misalignment
@@
  • Semantic variables $\mathbf{s}\in\mathcal{S}\subseteq\mathbb{R}^{n_{s}}$: Latent variables capturing the semantic content of the data, i.e., information that is interpretable or describable through human knowledge (e.g., object shape, color). We denote the index set of semantic variables as $\mathbb{I}_{\mathbf{s}}:=[n_{s}]$ [^3] Throughout this paper, we use the notation $[d]$ to represent the set $\{1,\dots,d\}$ for any integer $d>1$. for future reference.
  • Image-specific variables $\mathbf{m}_{x}\in\mathcal{M}_{x}$  [^4] For simplicity of notation, we omit the dimensions of certain variables throughout this work, such as $\mathbf{m}_{x}$.: Latent variables capturing image-specific, non-semantic factors that are independent of semantic variables $\mathbf{s}$ (e.g., camera noise or background artifacts).
  • Text-specific variables $\mathbf{m}_{t}\in\mathcal{M}_{t}$: Latent variables capturing text-specific, non-semantic factors that are independent of both semantic variables $\mathbf{s}$ and image-specific variables $\mathbf{m}_{x}$ (e.g., grammar).
A key advantage of the proposed latent space structure is its flexibility and applicability to real-world scenarios. To this end, we depart from prior works that impose somewhat restrictive assumptions on the latent structure. Specifically, unlike approaches that enforce certain fixed graphical structures among semantic variables (e.g., assuming content causally determines style) [12], or methods based on nonlinear ICA that assume complete (or conditional on an auxiliary variable) independence among latent variables [26, 61], we allow for arbitrary dependency structures among semantic variables $\mathbf{s}$, since the true latent graph structure among semantic variables is often unknown in practice.
@@
An image $\mathbf{x}\in\mathcal{X}$ is generated from latent variables $\mathbf{z}_{x}=(\mathbf{s},\mathbf{m}_{x})$ via a diffeomorphism (i.e., a bijection with smooth inverse function) $\mathbf{g}_{x}:\mathcal{S}\times\mathcal{M}_{x}\to\mathcal{X}$:
@@
Unlike other multimodal data (e.g., camera-LiDAR [35]) acquired via sensors, given an image, its corresponding text inherently exhibits flexibility in semantic richness and may also include distortions introduced by humans or captioning models [34, 73], leading to cross-modal misalignment. Here, we formalize the root of such misalignment by introducing two types of biases: selection bias and perturbation bias. Accordingly, before formulating the text generation process, we first provide definitions of these two types of biases.
@@
Let $\mathcal{P}_{+}(\mathbb{I}_{\mathbf{s}})$ denote the set of all non-empty subsets [^5] Without loss of generality, we fix a graded lexicographic order over $\mathcal{P}_{+}(\mathbb{I}_{\mathbf{s}})$. See Clar. B.2 for a clarification. of the index set $\mathbb{I}_{\mathbf{s}}$, defined as $\mathcal{P}_{+}(\mathbb{I}_{\mathbf{s}}):=\mathcal{P}(\mathbb{I}_{\mathbf{s}})\setminus\{\emptyset\}$, where $\mathcal{P}(\cdot)$ denotes the power set. The selection bias $\theta$ is defined as an integer index in the range $[2^{n_{s}}-1]$, corresponding to a specific non-empty semantic subset $\mathbb{I}_{\theta}\in\mathcal{P}_{+}(\mathbb{I}_{\mathbf{s}})$. The complement $\mathbb{I}_{\theta}^{c}=\mathbb{I}_{\mathbf{s}}\setminus\mathbb{I}_{\theta}$ denotes the omitted semantic subset.
@@
Let $\mathcal{P}_{\text{proper}}(\mathbb{I}_{\theta}):=\mathcal{P}(\mathbb{I}_{\theta})\setminus\{\mathbb{I}_{\theta}\}$ denote the set of all proper subsets [^6] Again, we fix a graded lexicographic order over the proper subsets in $\mathcal{P}_{\text{proper}}(\mathbb{I}_{\theta})$ throughout the paper. of the selected index subset $\mathbb{I}_{\theta}$. The perturbation bias $\rho$ is defined as an integer index in the range $\rho\in[2^{|\mathbb{I}_{\theta}|}-1],$ corresponding to a unique subset $\mathbb{I}_{\rho}\in\mathcal{P}_{\text{proper}}(\mathbb{I}_{\theta})$ subject to perturbation. The complement $\mathbb{I}_{\rho}^{c}:=\mathbb{I}_{\theta}\setminus\mathbb{I}_{\rho}$ denotes the semantic subset that remains unperturbed across modalities.
@@
where $p_{\mathbf{m}_{t}}$ denotes the prior distribution over the latent subspace $\mathcal{M}_{t}$, and $\mathbf{g}_{t^{\scriptscriptstyle(\theta)}}$ is the diffeomorphic mapping specified by the selection bias $\theta$. This formulation explicitly characterizes misalignment in text generation arises through perturbations within selected semantic dimensions.